Ask your own question, for FREE!
Mathematics 8 Online
OpenStudy (anonymous):

Krissy has 3 arrows. The probability of her hitting the target is 2/5. What is the probability that she will get at least one arrow on the target?

OpenStudy (mathmate):

At least one means she does not miss all, which is easier to calculate. Probability of missing all three is P(0)=(3/5)^3 so P(1+)=1-P(0)

OpenStudy (anonymous):

What would the answer be in percent?

OpenStudy (mathmate):

P(0)=(3/5)^3= 0.216 so P(1+)=1-P(0)=0.784

OpenStudy (anonymous):

Thank you!

OpenStudy (mathmate):

You're welcome!

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
2 days ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
2 days ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!